Ceph/CVMFS/Filer Service Meeting

Europe/Zurich
600-R-001 (CERN)

600-R-001

CERN

15
Show room on map
    • 14:00 14:15
      Visitor Talk: CephFS Benchmarking for ATLAS 15m
      Speaker: Adam Abed Abud (Universita and INFN (IT))
    • 14:15 14:20
      CVMFS 5m
      Speaker: Enrico Bocchi (CERN)

      Enrico:

      • WIP atlas-nightlies on S3
        • Issues with CC7 (they have never used ayum with CC7)
        • Extension of the Manila share for $HOME
        • Long publication time due to many 'HEAD' requests against S3
      • New set of squids for Atlas repos
      • Extension of atlas.cern.ch volume (4.5TB)
      • Memory allocation problem with cvmfs-server-2.5.2 (`_CVMFS_SERVER_PIPELINE_MB=64`)
      • Migrated {belle, boss, geant4, glast}.cern.ch to more powerful VM
    • 14:20 14:25
      Ceph Rota Report 5m

      Julien:

       - 3 disk replacements

       - meeting with the repair service this week on improving procedures?

    • 14:25 14:30
      Ceph Upstream News 5m

      Releases, Tickets, Testing, Board, ...

      Speaker: Dan van der Ster (CERN)

      From Dan

      • Mimic 13.2.5 was released. This *should* be the good one to start our big upgrades. Will prepare some jira's to test and start upgrades.
      • Nautilus stable nearly released. The final release notes are a good read: https://github.com/ceph/ceph/pull/27019
        Interesting to note is that RH and SUSE enterprise Ceph will update directly to Nautilus from Luminous.
    • 14:30 14:35
      Ceph Backends 5m

      Upgrades, capacity changes, rebalancing, ...

      Speaker: Theofilos Mouratidis (National and Kapodistrian University of Athens (GR))

      ceph/flax: 2nd host, first 6+1 disks formatted to bluestore, now backfilling

      ceph-scripts: avoid usage of `ceph-volume lvm zap` on `ceph-disk` formatted disks, because they start the osds and then they try to delete them. (i use ceph-disk zap now, it may need improvement because it creates a unusable gpt partition table)

      Didn't see any problems with restarting hosts with reformatted disks.
      They booted normally and started all processes.

       

       

      Dan

      • cephmon0 being deleted next week. 
        Need to identify new mons for beesly. Phys/VM is being debated. (Bootstrapping issue)
    • 14:35 14:40
      S3 5m

      Ops, Use-cases (backup, DB), ...

      Speakers: Julien Collet (CERN), Roberto Valverde Cameselle (Universidad de Oviedo (ES))

      Roberto (Backup):

      •  Working on a restic puppet module to easily install restic and schedule a backup to S3.

       

      Julien:

      • s3.cern.ch upgraded to the latest Ceph v12.2.11.
      • setting up and testing of cosbench to measure current S3 performance

      Herve (by Dan)

      • OTG0048860: S3 had only one IP in s3.cern.ch, following an updated lbclient rpm and our incompatible configuration. https://gitlab.cern.ch/ai/it-puppet-hostgroup-ceph/commit/5209d601b1eaccb096f7363e6a33423db7ba238d

      Dan

      • We need to prepare the configuration for multi-region S3.
        Roberto tested multi-region previously.. any documentation?
        We could create a tiny VM cluster now (with exact same zone config as gabe cluster), then create a 2nd tiny VM cluster and work out the procedure to add the 2nd zone.
    • 14:40 14:45
      Block Storage 5m

      OpenStack Cinder, Beesly, Wigner Decommissioning, ...

      Speaker: Theofilos Mouratidis (National and Kapodistrian University of Athens (GR))

      Dan

      • beesly balancing of RA21 is getting there: currently 66% used (from peak of 85% used roughly 1 month ago)
    • 14:45 14:50
      CephFS/FILER 5m
      Speakers: Alberto Chiusole (Universita e INFN Trieste (IT)), Dan van der Ster (CERN)

      Dan

      • OpenShift filer needed a hard reboot late Friday evening: OTG0048862
      • HPC 1m
        Speaker: Alberto Chiusole (Universita e INFN Trieste (IT))

        Running an HPC application (RegCM, Regional Climate Model - https://gforge.ictp.it/gf/project/regcm/) on CEPH /bescratch (kernel mounted), on different # procs.

        Strange "patterns": investigation with IOR in different ways of writing (MPI-IO, HDF5, NetCDF, etc)

    • 14:50 14:55
      HyperConverged 5m
      Speakers: Jose Castro Leon (CERN), Julien Collet (CERN), Roberto Valverde Cameselle (Universidad de Oviedo (ES))

      Jose:

      • Mail and DB guys are testing the hyperconverged setup
      • We may need to test enabling some tunables that are currently not there
      • Look at the impact of the cache configuration on the Client side
    • 14:55 15:00
      Monitoring 5m

      Julien:

      • Prophetstore set up on a couple of erin hosts: presumably one disk is to fail this or next week.
      • Willing to increase the free trial period
      • Larger-scale tests? (currently 3 erin hosts)

      Roberto:

      • Updated today to thanos v0.3.2, should improve memory usage by the compactor.
      • Some gaps observed in long-term metrics, need to check.

      Dan:

      • New tickets needing a volunteer:
        • Create KPI/SLA dashboards: https://its.cern.ch/jira/browse/CEPH-679
        • rationalize the ceph-health-cron: https://its.cern.ch/jira/browse/CEPH-680
    • 15:00 15:05
      AOB 5m

      Julien

      • Guys from IBM sent a ping about the LinuxOne machine

      RAL workshop:

      • Was useful -- they have a 5000 OSD cluster and enjoy some of the issues we also see (balancing, how to handle inconsistent PGs)
      • Slides available here: https://indico.cern.ch/event/803456/overview
      • Some dev ideas coming out:
        • better handling for weak writes (which leads to inconsistent PG): repair the PG immediately, and *overwrite* the object in place so that it may fix the PendingSector.
        • add pg-upmap-items-force (and rm-pg-upmap-items-force) which do not do any crush rule validation. This would be useful when migrating to new crush topologies. 
      •  
Your browser is out of date!

Update your browser to view this website correctly. Update my browser now

×